log2(x) + log2(8) = log2(64 − x) + 3
Solve for x

Answer by stanbon(75887)

log2(x) + log2(8) = log2(64 − x) + 3
Solve for x
--------------------
log2(x) + 3 = log2(64-x) = 3
----
log2(x) = log2(64-x)
---
x = 64-x
2x = 64
x = 32
===============
Cheers,
Stan H.
======================